# Hardness-Dependent Optimized Quantum Adiabatic Schedules for Maximum Independent Sets

### v1

@@ II.3 Results
In Fig. 4, we compare the probability $P_{\text{MIS}}$ to reach the solution for 4 protocols: the piecewise linear and ACQC protocols defined in [27] which serve as benchmarks, and their generalized and hardness-dependent counterparts described in the previous sections. Note that $a=5.5$ for the benchmark protocols but $a=5.0$ in our study. The $\mathcal{HP}$-linear protocol yields the best results for all graphs and in particular for the hard ones, with a tenfold improvement over the benchmark protocol in this case. Detailed results for all graphs and protocols are illustrated in Fig. S8.
@@
As for the optimized parameters, we observe that the values of $P_{\text{MIS}}$ depend mostly on the hardness parameter and not on the order of the graphs. For instance, the success probabilities for the $\mathcal{HP}$-linear protocol range from 0.59 ($\mathcal{HP}=4.20$) to 0.99 ($\mathcal{HP}=0.46$) for $N=10$, and from 0.11 ($\mathcal{HP}=9.71$) to 0.98 ($\mathcal{HP}=0.52$) for $N=17$.
@@
Figure 4: Probability to reach a MIS depending on the hardness for the 500 graphs of the dataset. The piecewise linear (blue) and ACQC (orange) protocols serve as benchmarks. For hard graphs, the generalized $\mathcal{HP}$-ACQC (green) and $\mathcal{HP}$-linear (red) protocols developed in this study lead to an average success probability that is improved by approximately 150% and 1 000%, respectively.
III Experiments
@@
Based on thorough numerical simulations, we have shown in the previous section that hardness-dependent schedules lead to high probabilities to reach the MIS solution. Now, we extend these results to the hardware and to graphs that are not part of the dataset used in the optimization. First, we consider a small set of “toy graphs” that we can still simulate classically to ensure that the various protocols are correctly run in Aquila. Then, we apply the models to large graphs (137 vertices) to verify that the hardness parameter and not the order of the graph is indeed the most relevant property when solving the MIS problem.
@@
For all examples, we use 720 shots (repetitions of the adiabatic evolution) to get enough data for a statistical comparison of the methods. We choose this specific value because it is a multiple of 9, 12, and 16, which are the maximum number of copies of the small graphs ($3\times 3$ and $3\times 4$ lattices with $a=5$ and $a=5.5$) that fit and can be run simultaneously in the hardware. We consider only the shots in which the lattice is correctly filled by atoms. The typical probability of failing to occupy a site in Aquila is $\epsilon_{\text{fill}}=0.007$ [20, Sec. 1.5], and in practice we get $684\pm 7$ and $380\pm 14$ valid shots for the small and the large graphs, respectively.
III.1 Toy graphs
@@
In Fig. 5, we display the probability to reach the MIS solution for the 11 toy graphs defined in [28] and detailed in Sec. SIII.1. As in that article, we set $t_{\text{f}}=0.7$ so that the probabilities of success are distributed around one half. We observe a good agreement between numerical simulations and the corresponding results obtained on hardware, especially for the benchmark protocols. One possible reason for the optimized schedules to yield outcomes that are slightly lower than expected may be the choice of the larger lattice spacing, or the larger gradient in $\Omega$; many other sources of error could explain these results, see [20, Sec. 1.5] for an exhaustive list. Nevertheless, the $\mathcal{HP}$-linear protocol still performs better on the hardware than the linear and CD benchmarks, with a threefold and a twofold mean improvement, respectively.
@@
Figure 5: Probability to reach the MIS solution for the 11 toy graphs. The rectangle bars are numerical results and the error bars are experimental results (95% confidence intervals). The ranking of the 4 studied protocols is the same as in Fig. 4, with $\mathcal{HP}\in[2,3]$ here. The black horizontal line represents the average success probability obtained by the Bayesian optimization in [28, Fig. 9], which performs better than the benchmarks (blue and orange) but not as well as the hardness-dependent protocols (green and red).
III.2 Large graphs
@@
We now turn to large graphs, which are of much higher interest in real applications. We consider three of the four graphs studied in [28, Fig. 10], which consist of 137 vertices and whose hardness parameters span three orders of magnitude ($\mathcal{HP}\approx$ 14, 125, and 1435, with the second one referred as “95” in the mentioned article). In addition, we generate one very easy graph of the same size to span yet another order of magnitude ($\mathcal{HP}\approx$ 1.4), see Fig. 6. The hardness parameter of these large graphs is efficiently computed by advanced classical algorithms, namely by the GenericTensorNetworks Julia library [38, 39].
@@
Since the MIS solutions and consequently the hardness parameters are not known a priori, we apply our optimized linear schedules in the limit $\mathcal{HP}\to\infty$. Moreover, we set $t_{\text{f}}=t_{\text{max}}=4$ so that the probability to find a MIS solution is as high as possible. Finally, we post-process all experimental data such that the measured outcomes become valid maximal independent sets. To this end, we use the greedy procedure described in [18, Sec. S2.3]: first, we remove any of two neighboring vertices until the sets become independent. Second, we add vertices to the free sites with no neighbor until the sets are maximal. The same procedure applied to an empty configuration is used as a classical benchmark, which highlights the relative gain obtained with the quantum protocols [20, Sec. 6].
@@
Results are summarized in Fig. 6: the MIS solution is never found by the classical greedy algorithm and is obtained by the standard linear protocol only for the easiest graph. On the contrary, the $\mathcal{HP}$-linear protocol leads to the solution for all but the hardest graph, which indicates a clear improvement over the existing protocols.
IV Discussion and outlook
@@
In this work, we have shown how to carefully design the adiabatic schedules of the Rydberg Hamiltonian to efficiently solve the MIS problem on unit-disk graphs.
@@
First, a statistical investigation of the conditions for which the final quantum state encodes the MIS solution has revealed that the final detuning should lie in a narrow interval. To our knowledge, these are the tightest bounds ever mentioned. We believe that choosing a final detuning in that interval would improve the results of many other studies, in terms of success probability or number of optimization steps. In fact, applying adiabatic protocols in which the final ground state does not correspond to the sought solution can only lead to poorly performing algorithms.
@@
Second, the individual optimization of numerous small graphs has shown that the optimum parameters of the schedules depend mostly on the hardness parameter and only marginally on the size of the graph. Through extensive classical simulations of small systems, we have generated adiabatic schedules that perform extremely well also when applied to much larger graphs; only the hardest instances are not solved by our protocol. Obviously, improvements in the hardware would greatly help in this regard, especially if it permits longer coherent evolution times. Since the quantum technology evolves rapidly, this could be the case soon. However, because of the combinatorial nature of the hardness parameter, very hard MIS problems exist only in sufficiently large graphs, so that our numerical study should be extended to larger graphs to address this issue. We believe that improved schedules could be easily found by considering slightly larger systems, for instance up to about 25 vertices, or even larger using tensor network techniques whenever appropriate [40].
@@
We have focused our work on two families of schedules, namely piecewise linear and smooth counterdiabatic functions. The phase of the Rabi drive is non constant only in the latter case, and we have developed a generalized version of the ACQC protocol [27] to make it specific to each graph. However, when correctly parametrized, the simpler $\mathcal{HP}$-linear protocol performs better in almost all numerical simulations. The difference is even more pronounced in experiments, as the schedules have to be described in terms of piecewise linear or constant functions. We have used a maximum number of 6 free parameters in the optimization, but more general functions may be investigated, as the method described in this work applies to any kind of schedules. Moreover, it would be interesting to develop models of optimized schedules that do not depend only on the hardness parameter but also on other global properties of the graphs, e.g., the mean degree or the mean eccentricity of the vertices. Based on machine learning, such additional predictors could in fact improve the fitting of the schedules.
@@
We think that hybrid classical-quantum algorithms could take advantage of our protocols and, more generally, of precomputed optimized schedules as described in this study: instead of searching in the huge space of all parameters, one could consider lower-dimensional refined ansatzes, such as the $\mathcal{HP}$-parametrization of the schedules or the principal component of the residuals of the fitted parameters.
@@
Finally, we believe that the piecewise linear protocols developed in this study should serve as a new benchmark for any future adiabatic algorithm that aims to solve the MIS problem with Rydberg atoms. In particular, the most basic version of this protocol with 4 free parameters and in the limit of infinite hardness admits a very simple mathematical description, see Fig. 2, and it leads to relatively good results regardless of the specific graph to be solved.
@@
Figure 6: Experiments on large unit-disk graphs. Top: four graphs with 137 vertices each and whose hardness parameters span four orders of magnitude; the three hardest graphs are taken from [28]. In shades of blue, the probability $\braket{\hat{n}_{i}}$ to measure the $i$-th atom in the Rydberg state |$\ket{1}$ at the end of the $\mathcal{HP}$-linear protocol. Due to the high degeneracy of the MIS, some vertices are equally measured in the ground or excited state. Bottom: experimental results after post-processing to find an independent set of given size, with the MIS solution at the rightmost bin; error bars represent the 95% confidence intervals. The classical greedy algorithm (gray) fails to find the solution in all cases. The benchmark linear protocol (blue) reaches the MIS in the easiest case but fails for any harder instance. Finally, the $\mathcal{HP}$-linear protocol is able to find the MIS except for the hardest graph; the corresponding values of $P_{\text{MIS}}$ are indicated in the subtitles. Results for the two CD protocols (not shown) lie between those of the two linear protocols.
Supplemental Material
SI Ground state encoding of MIS
@@
Given a simple undirected graph $G=(V,E)$ with vertex set $V$ and edges set $E$, a subset $S\subseteq V$ is called independent if no pair of vertices in $S$ are connected by an edge in $E$. In the maximum independent set (MIS) problem, we need to find the independent set of largest cardinality. Denoting by the binary variable $x_{i}$ the presence of the vertex $i=1,\ldots,N$ in $S$, it is proven that a MIS solution minimizes the cost function [42, 43]
$\mathcal{C}(x_{1},\ldots,x_{N})=-\sum_{i\in V}\Delta\,x_{i}+\sum_{ij\in E}V_{ij}x_{i}x_{j}$ (S1)
@@
if $0<\Delta<V_{ij}$ for all $ij\in E$. In our setting, $G$ is a unit-disk graph that connects the nearest and next-nearest neighbors of a square lattice. Neglecting the interactions between farther atoms and for a non-planar graph (i.e., when crossing diagonals belong to it), the MIS condition becomes:
$0<\Delta_{\text{f}}<\frac{C_{6}}{(\sqrt{2}\,a)^{6}}=\frac{V_{0}}{8},$ (S2)
where $a$ is the lattice spacing (considered as a fixed parameter), $V_{0}=C_{6}/a^{6}$ is the interaction strength between adjacent atoms, and $\Delta_{\text{f}}=\Delta(t_{\text{f}})$ is the detuning at the final evolution time $t_{\text{f}}$. Obviously, for the Rydberg Hamiltonian to encode the MIS cost function at $t_{\text{f}}$, we also need $\Omega(t_{\text{f}})=0$, see Eq. (1).
@@
In several investigations, it is implicitly assumed that the inequalities (S2) hold for all unit-disk graphs due to the sharp Rydberg blockade transition. However, in what follows we show that the condition $\Delta_{\text{f}}>0$ is necessary but not sufficient.
SI.1 Nonexistence of a lower bound on the detuning
@@
The Van der Waals interaction decays quickly with the distance between atoms, but it cannot be neglected for distances close to the blockade radius. As an example, consider the graph that consists of three atoms in a row with a distance $2a$ between the extremities, see Fig. S1(a). In this case, the MIS solution is $(1,0,1)$, which corresponds to an energy $-2\Delta_{\text{f}}+V_{0}/64$. By listing the energy of all other configurations, we find the following conditions on $\Delta_{\text{f}}$ such that |$\ket{101}$ is the ground state of $H(t_{\text{f}})$:
$V0/64<\Delta_{\text{f}}<2\,V_{0}.$ (S3)
The upper bound is of course already satisfied in Eq. (S2) as it is a sufficient condition, but the value $V_{0}/64$ should be considered as a strict lower bound for any protocol; remark that this bound is mentioned in [20].
Figure S1: Examples of unit-disk graphs to study the lower bound on the final detuning. The MIS solution (filled circles) has a lower energy than the suboptimal configuration (orange circles) iff $\Delta_{\text{f}}>\Delta_{\text{LB}}$. (a) Reference graph used in Eq. (S3). (b) Graph with $N=3m+1$ vertices forming $m$ alternating upwards and downwards triangles. For any $\Delta_{\text{f}}>0$, there exists a value of $m$ such that the suboptimal configuration is energetically favorable. Finally, two non-planar graphs with very different lower bounds: (c) $\Delta_{\text{LB}}\approx V_{0}/1017$ ($\mathcal{HP}\approx 0.42$), and (d) $\Delta_{\text{LB}}\approx V_{0}/11.49$ ($\mathcal{HP}=15$).
@@
The question is then to determine if there exists, for unit-disk graphs in a square lattice, a finite lower bound on the detuning to ensure the equivalence between the classical MIS solution and the ground state of the quantum Hamiltonian. Unfortunately, the answer is negative. In fact, consider the unit-disk graph in Fig. S1(b), and the following two sets of vertices:
@@
    the $m$ vertices located at the tip of the alternating triangles (orange circles)
@@
$E_{0}$ $=-(m+1)\,\Delta_{\text{f}}+\frac{V_{0}}{64}\sum_{i=1}^{m}\frac{m+1-i}{i^{6}},$ (S4a)
$E_{1}$ $=-m\,\Delta_{\text{f}}+\frac{V_{0}}{64}\sum_{i=1}^{\lfloor m/2\rfloor}\left(\frac{m+1-2i}{(4i^{2}-4i+2)^{3}}+\frac{m-2i}{(2i)^{6}}\right).$ (S4b)
@@
$\frac{V_{0}}{64}\left(\Sigma_{0}-\Sigma_{1}\right)<\Delta_{\text{f}},$ (S5)
@@
where $\Sigma_{0,1}$ stands for the sums in the r.h.s. of Eq. (S4). However, one can check that the term $\Sigma_{0}-\Sigma_{1}$ is always positive and becomes arbitrarily large for increasing $m$. In particular, this difference is larger than 8 for $m\geq 9$, leading to a contradiction with Eq. (S2).
@@
In conclusion, there does not exist a value of the final detuning such that the Rydberg Hamiltonian is ensured to encode the MIS solution of all unit-disk graphs.
SI.2 Statistical analysis of the bounds on the detuning

### v2

@@ 2.3 Results
In Fig. 6, we compare the probability $P_{\text{MIS}}$ to reach the solution for 4 protocols: the piecewise linear and ACQC protocols defined in [28] which serve as benchmarks, and their generalized and hardness-dependent counterparts described in the previous sections. Note that $a=5.5$ for the benchmark protocols but $a=5$ in our study. The $\mathcal{HP}$-linear protocol (with 6 degrees of freedom) yields the best results for all graphs, and in particular for the hard ones with a tenfold improvement over the benchmark.
@@
Graph-to-graph comparisons are illustrated in Fig. 7. As for the optimized schedule parameters, we observe that the values of $P_{\text{MIS}}$ depend mostly on the hardness parameter and not on the size of the graphs. For instance, the success probabilities for the $\mathcal{HP}$-linear protocol range from 0.59 ($\mathcal{HP}=4.20$) to 0.99 ($\mathcal{HP}=0.46$) for $N=10$, and from 0.11 ($\mathcal{HP}=9.71$) to 0.98 ($\mathcal{HP}=0.52$) for $N=17$.
@@
Figure 6: Box plot of the probability to reach the MIS depending on the hardness of the 500 graphs of the dataset. The piecewise linear (blue) and ACQC (orange) protocols serve as benchmarks. For hard graphs, the generalized $\mathcal{HP}$-ACQC (green) and $\mathcal{HP}$-linear (red) protocols developed in this study lead to an average success probability that is improved by approximately 150% and 1 000%, respectively.
Figure 7: Graph-to-graph comparisons of $P_{\text{MIS}}$. Left: the $\mathcal{HP}$-linear protocol developed in this study leads to a 3.6-fold mean improvement on the linear protocol described in [28]. Right: except for rare instances, the hardness-dependent piecewise linear protocol always beats its ACQC counterpart. Note that $P_{\text{MIS}}$ depends strongly on $\mathcal{HP}$ but not on $N$ (upward or downward triangle), with low sucess probabilities for the hard graphs (maroon) and nearly perfect results for the easy ones (green).
3 Experiments
@@
Based on thorough numerical simulations, we have shown in the previous section that hardness-dependent schedules lead to high probabilities to reach the MIS solution. Now, we extend these results to the hardware and to graphs that are not part of the dataset used in the optimization. First, we consider a small set of “toy graphs” that we can still simulate classically to ensure that the various protocols are correctly run in Aquila. Then, we apply the models to large graphs made up of 137 vertices to verify that the hardness parameter and not the size of the graph is indeed the most relevant property when solving the MIS problem.
@@
For all examples, we use 720 shots (repetitions of the adiabatic evolution) to get enough data for a statistical comparison of the methods. We choose this specific value because it is a multiple of 9, 12, and 16, which are the maximum number of copies of the small graphs ($3\times 3$ and $3\times 4$ lattices with $a=5$ and $a=5.5$) that fit and can be run simultaneously in the hardware. We consider only the shots in which the lattice is correctly filled by atoms. The typical probability of failing to occupy a site in Aquila is $\epsilon_{\text{fill}}=0.007$ [21, Sec. 1.5], and in practice we get $684\pm 7$ and $380\pm 14$ valid shots for the small and the large graphs, respectively.
3.1 Toy graphs
@@
In Fig. 8, we display the probability to reach the MIS solution for the 11 toy graphs defined in [29]. As in that article, we set $t_{\text{f}}=0.7$ so that the probabilities of success are distributed around 50%. We observe a good agreement between numerical simulations and the corresponding results obtained on hardware, especially for the benchmark protocols. A possible reason for the optimized schedules to yield outcomes that are slightly lower than expected may be the choice of a smaller lattice spacing or the larger gradient of $\Omega$ at the end of the evolution; many other sources of error could explain these differences, see [21, Sec. 1.5] for an exhaustive list. Nevertheless, the $\mathcal{HP}$-linear protocol still performs better on the hardware than the linear and CD benchmarks, with a threefold and a twofold mean improvement, respectively.
@@
Figure 8: Probability to reach the MIS solution for the 11 toy graphs. The rectangle bars are numerical results and the error bars are experimental results (95% confidence intervals). The ranking of the 4 studied protocols is the same as in Fig. 6, with $\mathcal{HP}\in[2,3]$ here. The black horizontal line represents the average success probability obtained by the Bayesian optimization in [29, Fig. 9], which performs better than the benchmarks (blue and orange) but not as well as the hardness-dependent protocols (green and red).
3.2 Large graphs
@@
We now turn to large graphs, which are of much higher interest in real applications. We consider three of the four graphs studied in [29, Fig. 10], which consist of 137 vertices and whose hardness parameters extend over three orders of magnitude ($\mathcal{HP}\approx$ 14, 126, and 1435). In addition, we generate one very easy graph of the same size to span yet another order of magnitude ($\mathcal{HP}\approx$ 1.4), see Fig. 9. The hardness parameter of these large graphs is efficiently computed by advanced classical algorithms, namely by the GenericTensorNetworks Julia library [37, 38].
@@
Since the MIS solutions and consequently the hardness parameters are not known a priori, we apply our optimized linear schedules in the limit $\mathcal{HP}\to\infty$. Moreover, we set $t_{\text{f}}=t_{\text{max}}=4$ so that the probability to find a MIS solution is as high as possible. Finally, we post-process all experimental data such that the measured outcomes become valid maximal independent sets. To this end, we use the greedy procedure described in [19, Sec. S2.3]: first, we remove any of two neighboring vertices until the sets become independent. Second, we include a vertex at any free site with no neighbor until the sets are maximal. The same procedure applied to an empty configuration is used as a classical benchmark, which highlights the relative gain obtained with the quantum protocols [21, Sec. 6].
@@
The results are summarized in Fig. 9: the MIS solution is never found by the classical greedy algorithm and is obtained by the standard linear protocol only for the easiest graph. On the contrary, the $\mathcal{HP}$-linear protocol leads to the solution for all but the hardest graph, which indicates a clear improvement over the existing protocols.
4 Discussion and outlook
@@
In this work, we have shown how to design the adiabatic schedules of the Rydberg Hamiltonian to efficiently solve the MIS problem on unit disk graphs.
@@
First, a statistical investigation of the conditions for which the final quantum state encodes the MIS solution has revealed that the final detuning should lie in a narrow interval. To our knowledge, these are the tightest bounds ever mentioned. We believe that choosing a final detuning in that interval would improve the results of many other studies, either in terms of success probability or number of optimization steps. In fact, applying adiabatic protocols in which the final ground state does not correspond to the sought solution can only lead to poorly performing algorithms.
@@
Second, the individual optimization of numerous small graphs has shown that the optimum parameters of the schedules depend mostly on the hardness parameter and only marginally on the size of the graph. Through extensive classical simulations of small systems, we have generated adiabatic schedules that perform extremely well also when applied to much larger graphs; only the hardest instances are not solved by our protocol. Obviously, improvements in the hardware would greatly help in this regard, especially if it permits a longer coherent evolution time. Since the quantum technology evolves rapidly, this could be the case soon. However, because of the combinatorial nature of the hardness parameter, very hard MIS problems exist only in sufficiently large graphs, so that our numerical study should be extended to larger graphs to address this issue. We believe that improved schedules could be easily found by considering slightly larger systems, for instance up to about 25 vertices, or even larger ones using tensor network techniques for the simulations whenever appropriate [39].
@@
We have focused our work on two families of schedules, namely piecewise linear and smooth counterdiabatic functions. The phase of the Rabi drive is nonconstant only in the latter case, and we have developed a generalized version of the ACQC protocol [28] to make it specific to each graph. However, when correctly parametrized, the somehow simpler $\mathcal{HP}$-linear protocol performs better in almost all numerical simulations. This difference is even more pronounced in experiments, as the smooth CD schedules have to be converted into piecewise linear or constant functions. We have used a maximum number of 6 free parameters in the optimization, but more general functions may be investigated since the method described in this work applies to any kind of schedules. Moreover, it would be interesting to develop models of optimized schedules that do not depend only on the hardness parameter but also on other global properties of the graphs, e.g., the mean degree or the mean eccentricity of the vertices. Based on machine learning, such additional predictors could in fact improve the fitting of the schedule parameters.
@@
We think that hybrid classical-quantum algorithms could take advantage of our protocols and, more generally, of precomputed optimized schedules: instead of searching in the huge space of all parameters, one could consider lower-dimensional meaningful ansatzes, such as the $\mathcal{HP}$-parametrization of the schedules or the principal component of the residuals of the fitted parameters.
@@
Finally, we believe that the piecewise linear protocols developed in this study should serve as a new benchmark for any future adiabatic algorithm that aims to solve the MIS problem with Rydberg atoms. In particular, the most basic version of this protocol with only 4 free parameters and in the limit of infinite hardness admits a very simple mathematical description, see Fig. 2, and it leads to relatively good results regardless of the specific graph to be solved.
@@
Figure 9: Experiments on large unit disk graphs. Top: four graphs with 137 vertices each and whose hardness parameters span four orders of magnitude; the three hardest graphs are taken from [29]. In shades of blue, the probability $\braket{\hat{n}_{i}}$ to measure the $i$-th atom in the Rydberg state |$\ket{1}$ at the end of the $\mathcal{HP}$-linear protocol. Due to the high degeneracy of the MIS, some vertices are equally measured in the ground or excited state. Bottom: experimental results (after a minimal post-processing) to find an independent set of a given size, with the MIS solution at the rightmost bin. The error bars represent the 95% confidence intervals. The classical greedy algorithm (gray) fails to find the solution in all cases. The benchmark linear protocol (blue) reaches the MIS in the easiest case but fails for any harder instance. Finally, the $\mathcal{HP}$-linear protocol is able to find the MIS except for the hardest graph; the corresponding values of $P_{\text{MIS}}$ are indicated in the subtitles. The success probabilities for the two CD protocols (not shown) lie between those of the two linear protocols.
[…]
Appendix A Ground state encoding of the maximum independent set problem
@@
Given a simple undirected graph $\mathcal{G}=(V,E)$ with vertex set $V$ and edges set $E$, a subset $S\subseteq V$ is called independent if no pair of vertices in $S$ are connected by an edge in $E$. In the maximum independent set (MIS) problem, we want to find the independent set of largest cardinality. Denoting by $x_{i}\in\{0,1\}$ the absence or the presence of the vertex $i=1,\ldots,N$ in $S$, it is proven that the MIS solution minimizes the cost function [41, 42]
$\mathcal{C}(x_{1},\ldots,x_{N})=-\sum_{i\in V}\Delta\,x_{i}+\sum_{ij\in E}V_{ij}\,x_{i}x_{j}$ (9)
@@
if $0<\Delta<V_{ij}$ for all $ij\in E$. In analog quantum computing with arrays of Rydberg atoms, $\mathcal{G}$ is a unit disk graph that connects the nearest and next-nearest neighbors of a square lattice. Neglecting the interactions between farther atoms and for a non-planar graph (i.e., when crossing diagonals belong to it), the MIS condition becomes:
$0<\Delta_{\text{f}}<\frac{C_{6}}{(\sqrt{2}\,a)^{6}}=\frac{V_{0}}{8},$ (10)
@@
where $a$ is the lattice spacing, $V_{0}=C_{6}/a^{6}$ is the interaction strength between adjacent atoms, and $\Delta_{\text{f}}=\Delta(t_{\text{f}})$ is the detuning at the final evolution time $t_{\text{f}}$. In several investigations, it is implicitly assumed that the inequalities (10) hold for all unit disk graphs due to the sharp Rydberg blockade transition. However, in what follows we show that the condition $0<\Delta_{\text{f}}$ is necessary but not sufficient.
A.1 Nonexistence of a lower bound on the detuning
@@
The Van der Waals interaction decays quickly with the distance between atoms, but it cannot be neglected for distances close to the blockade radius. As an example, consider the graph that consists of three atoms in a row with a distance $2a$ between its extremities, see Fig. 10(a). In this case, the MIS solution is $(1,0,1)$, which corresponds to an energy $-2\Delta_{\text{f}}+V_{0}/64$. By listing the energy of all other configurations, we find the following conditions on $\Delta_{\text{f}}$ such that |$\ket{101}$ is indeed the ground state of the final Rydberg Hamiltonian:
$V_{0}/64<\Delta_{\text{f}}<2\,V_{0}.$ (11)
@@
The upper bound is of course already satisfied in Eq. (10) as it is a sufficient condition, but the value $V_{0}/64$ should be considered as a strict lower bound for any protocol; remark that this bound is mentioned in [21]. The question is then to determine if there exists, for unit disk graphs in a square lattice, a finite lower bound on the detuning to ensure the equivalence between the classical MIS solution and the ground state of the quantum Hamiltonian. Unfortunately, the answer is negative. In fact, consider the unit disk graph in Fig. 10(b), and the following two sets of vertices:
@@
    the $m$ vertices located at the tip of the alternating upward and downward triangles (orange circles)
@@
$E_{0}$ $=-(m+1)\,\Delta_{\text{f}}+\frac{V_{0}}{64}\sum_{i=1}^{m}\frac{m+1-i}{i^{6}},$ (12a)
$E_{1}$ $=-m\,\Delta_{\text{f}}+\frac{V_{0}}{64}\sum_{i=1}^{\lfloor m/2\rfloor}\left(\frac{m+1-2i}{(4i^{2}-4i+2)^{3}}+\frac{m-2i}{(2i)^{6}}\right).$ (12b)
@@
$\frac{V_{0}}{64}\left(\Sigma_{0}-\Sigma_{1}\right)<\Delta_{\text{f}},$ (13)
@@
where $\Sigma_{0,1}$ stands for the sums in the r.h.s. of Eq. (12). However, one can check that the difference $\Sigma_{0}-\Sigma_{1}$ is always positive and becomes arbitrarily large for increasing $m$. In particular, this difference is larger than 8 for $m\geq 9$, leading to a contradiction with Eq. (10).
@@
In conclusion, there does not exist a value of the final detuning such that the Rydberg Hamiltonian is ensured to encode the MIS solution of all unit disk graphs.
Figure 10: Examples of unit disk graphs to study the lower bound on the final detuning. The MIS solution (filled circles) has a lower energy than the suboptimal configuration (orange circles) iff $\Delta_{\text{f}}>\Delta_{\text{LB}}$. (a) Reference graph used in Eq. (11). (b) Graph with $N=3m+1$ vertices forming $m$ alternating upward and downward triangles. For any $\Delta_{\text{f}}>0$, there exists a value of $m$ such that the suboptimal configuration is energetically favorable. Finally, two non-planar graphs with very different lower bounds: (c) $\Delta_{\text{LB}}\approx V_{0}/1017$ ($\mathcal{HP}\approx 0.42$), and (d) $\Delta_{\text{LB}}\approx V_{0}/11.49$ ($\mathcal{HP}=15$).
A.2 Statistical analysis of the bounds on the detuning